# Cheaper and more noise-resilient quantum state preparation using eigenvector continuation

### v1

@@ III.2 Scaling with system size
Although it is highly dependent on the nature of the system, the number of basis vectors required tends to increase polynomially with system size. In cases like the XY model, which exhibits different symmetry sectors in its ground states, the number of protected crossings is proportional to the number of lattice sites (see Fig. A4), and consequently the subspace requires a similarly scaling number of basis states. For other systems like XXZ model (see appendix D.2), where the ground state is doubly degenerate and has no protected crossings in the ground state, the growth is still polynomial but slower. We demonstrate this numerically using truncated ASP and EC by calculating the RMS error as we increase the number of subspace basis vectors $k_{p}$ for different system sizes of the XY and XXZ models. We use the same parameters for the XY model as those in Fig. 2; for the XXZ model, we set $J=+1$, $B_{Z}=0.2$ and vary $J_{Z}\in[0,3]$ using truncated ASP with parameters $dt=0.1$ and $T_{\mathrm{max}}=1.2$. Note that for the XXZ model ASP is easier to implement due to the absence of protected crossings; the $B_{Z}$ field is introduced to break the (two-fold) degeneracy in the ground state.
@@
The low-energy subspace typically only contains a finite number of unique states. As we increase the number of training points, it is possible that we overspan the low-energy space, making the $\mathbb{S}$ matrix non-invertible. Thus, we only calculate errors for $k_{p}\leq(N_{XY}+2)$ for the XY model, and $k_{p}\leq N_{XXZ}$ for the XXZ model, which were empirically determined based on when the subspace basis vectors become degenerate.
@@
For both of these models, we observe that the number of basis vectors required for $(\Delta\varepsilon_{\mathrm{rel}})_{rms}<5\%$ scales linearly with the system size $N$, in stark contrast to the exponential scaling of the Hilbert space dimension. For the XY model, this is because the number of phase transitions (or level crossings) in the ground energy spectrum is proportional to $N$ (see Section D.1). For the XXZ model, since the ground state changes smoothly without any level crossings (Fig. A5), we required an even smaller number of training states, that again appears to scale linear with system size ($N$).
@@ IV Avoiding protected level crossings and vanishing energy gaps
One of the benefits of using EC as a subspace method is that the subspace basis can be generated using Hamiltonians at points in parameter space that are easier to handle than the desired target parameters where state preparation methods perform poorly. For instance, ASP fails when the spectrum exhibits protected level crossings. In the previous section, these were broken by a small $B_{X}$ field; however, it may be that the $B_{X}=0$ problem is of interest. ITE can also have problems, and in particular has issues with small energy gaps between the ground and excited states. In this section, we will illustrate how to ameliorate this issue using EC.
@@ IV.1 Getting around protected level crossings in ASP
Figure 4: Ground energy spectrum for XY model 5-site with parameters $\{J=1,B_{X}=0.0\}$ calculated using SD with subspace obtained from ASP in parameter space with $\{J=1,B_{X}=0.1$}
As discussed in Sec. III.0.2, using ASP to generate a ground state requires finding a gapped adiabatic path, that is without any protected crossings [21], which precludes its use for systems that exhibit level crossings. The 1D transverse field XY model is such a system: in the absence of an in-plane magnetic field — the spectrum shows a number of level crossings equal to the number of sites — and thus ASP is not an appropriate tool for finding the ground state. However, we can get around this issue by performing ASP for the XY model that does have a small in-plane field, and build a subspace using training states from this model.
In the Fig. 4, we perform ASP for a system with $B_{X}=0.1$, $dt=0.05$ and $T_{\mathrm{max}}=3.75$, and choose four equally spaced states to form a basis. Next, we use this basis to find the eigenstates of the target system with $B_{X}=0.0$. As is clear from the figure, this procedure readily produces the correct ground state energies for the system of interest.
@@ V Noise resilience
In practice, since the currently available quantum hardware is noisy, we require algorithms to be noise resilient in order to yield useful results on a quantum device [11, 12, 13]. We already have some “noise” in the subspace in the form of excited states, arising from truncated state preparation methods; however, quantum devices introduce hardware noise as well, requiring us to analyze how additional noise will affect the algorithm. In this section, we will demonstrate that the improvements obtained using EC survive the addition of hardware noise. We implement a noise model and simulate the effect of introducing noise in the ASP for a 5-site XY model Hamiltonian discussed previously in Eq.  III.1.
@@
While increasing the number of training vectors (or subspace basis states) can help to reduce the error, this eventually reaches a barrier. Adding more basis vectors increases the chances of over-spanning the subspace, making the condition number $(C_{n})$ of the overlap matrix very large, and yielding it non-invertible (see Eq. 2). This is one of the major problems with subspace methods ([29, 31, 63]). To address this issue, there are several classical post-processing techniques used to manipulate the overlap matrix, such as thresholding [39]. This procedure involves finding the singular value decomposition of the overlap matrix, followed by eliminating the singular values that are smaller than a chosen threshold ($\alpha$). The vectors corresponding to the singular values less than $\alpha$ are discarded from the overlap matrix and the projected Hamiltonian matrix. This technique can help in reducing the condition number but sometimes we can lose some relevant information in the process. Let us analyze the effect of noise on condition number and thresholding for our test case.
@@ VI Discussion
In this work, we have used previously proposed eigenvector continuation as a subspace method [44], where continuity of the ground state of the system Hamiltonians varied by some order parameter is explored. In contrast to previous work, where exact eigenstates were needed to generate the subspace, we apply it to the subspace generated by truncated time evolution techniques (ASP and ITE) to generate a subspace. These truncated algorithms incorporate a number of low-energy states, which can be disentangled using EC, thus gaining better convergence while using fewer resources. Although here we have used time-evolution methods for generating a low-energy subspace, other methods such as ADAPT-VQE [14] or QAOA [21] can be similarly improved.
@@
We used various spin Hamiltonians with different ground state energy spectrum features to showcase the method, including the XY model with and without transverse field whose ground state manifold spans orthogonal symmetry sectors, the XXZ 1D model with transverse field which has small energy gaps between ground state and the first excited state, and the XXZ Kagome 2D model which has phase transitions and ground state degeneracies. However, this method is not limited to spin systems; the eigenvector continuation formalism has been previously used for other systems in nuclear physics, and quantum chemistry among others [46, 48, 45], and the truncated state preparation method proposed in this work can also be used for these systems.
All the findings shown in this work were performed on a classical simulator. Even here, compared to the full-time-evolution methods or exact diagonalization approach to get the ground states, using EC with truncated time evolution can be much faster. The implementation of subspace methods on quantum hardware has been previously studied [44, 64, 65], and it involves quantum circuits for obtaining the matrix elements of the overlap matrix ($\mathbb{S}$) and projected Hamiltonian ($\mathbb{H}$) with the basis vectors. This analysis carries over to the truncated methods discussed here; truncated quantum state preparation algorithms circuits are used to get approximate ground states as subspace basis vectors. Because we use the same basis for all the target Hamiltonians, the overlap matrix elements need to be calculated once. Similarly, for the projected Hamiltonian matrix, usually we do not need to repeat the measurements for each target Hamiltonian, since all the Pauli operators of the Hamiltonian are measured only once per basis set and only the coefficients are changed to calculated the target Hamiltonian [44]. The total number of circuits required scales quadratically in the number of basis states $k_{p}$: $k_{p}(k_{p}-1)$ for the overlap matrix and $k_{p}(k_{p}+1)$ for each of the Pauli operator in the projected Hamiltonian due to the structure of the matrices.
@@ Acknowledgements
This work was supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency.
@@ D.1 XY model
Figure A3: XY 1D chain system energy spectrum for $J=1$, varying $B_{Z}$ and $B_{X}\in\{0.0,0.2\}$. Level crossing split can be observed in the plot on the right.
Without an X- field, due to the presence of magnetization symmetry sectors, the ground state is protected for a range of varying $B_{Z}$. For particular values of $B_{Z}$, the ground state changes to some excited state resulting in a point of degeneracy where both states cross called level crossing. The ground state before and after these points is completely orthogonal. For a $N$ site 1D chain system, for each spin flip, we will have a magnetization phase transition making the number of symmetry sectors in 1D transverse field XY model with varying transverse field from $B_{Z}\in(-\infty,+\infty)$ as $(N+1)$. For example we plot $N=8$ energy spectrum with $J=1$, $B_{X}=0.0$ and $B_{Z}\in(-3,3)$ in Fig. A4. For the plots in this paper, we consider $B_{Z}\in[0,B_{Z(max)})$. Still, the scaling of the number of symmetry sectors would be linear in $N$.
@@
Adding a uniform field in the X-direction of strength $B_{X}$ has two significant consequences. One is the breaking of level crossings, and the other is the mixing of the lowest eigenstates. Breaking of level crossing introduces an energy gap between the lowest energy states, which makes the application of ASP possible. After adding a $B_{X}$ field, the ground states are not completely orthogonal unlike one without the field. This makes the training states mixed as well.
@@ D.2 XXZ model 1D chain
Figure A6: Implementing EC with truncated ASP and ITE for a 1D chain XXZ model 5 site. (a) and (d) plots the ground energy spectrum, (b) and (e) plots corresponding ground state fidelity change.
In the Fig. A6, truncated ITE and ASP are used to get a basis using and for XXZ 1D chain Hamiltonians D.2, with $J=1$, $B_{Z}=0.2$, and $J_{Z}\in(0,3)$ is the varying parameter. For truncated ITE in (a) and (b), $d\tau=0.2$ and $\tau_{\mathrm{max}}=1.2$ starting from a Haar random state. And for ASP in (c) and (d), starting from a ground state at $J_{Z}=0$, $4$ time steps with $dt=0.1$ and $T_{\mathrm{max}}=0.4$ are used to get the basis.
Figure A7: Energy spectrum of a Kagome Lattice system with $N_{X}=N_{Y}=2$, $J=1.0$, $B_{Z}=0.2$ and varying $J_{Z}$ showing lowest $100$ energy levels

### v3

@@ III.2 Scaling with system size
Although it is highly dependent on the nature of the system, the number of basis vectors required tends to increase polynomially with system size. In cases like the XY model, which exhibits different symmetry sectors in its ground states, the number of protected crossings is proportional to the number of lattice sites (see Fig. A4), and consequently, the subspace requires a similarly scaling number of basis states. For other systems like XXZ model (see appendix D.2), where the ground state is doubly degenerate and has no protected crossings in the ground state, the growth is still polynomial but slower. We demonstrate this numerically using truncated ASP and EC by calculating the RMS error as we increase the number of subspace basis vectors $k_{p}$ for different system sizes of the XY and XXZ models. We use the same parameters for the XY model as those in Fig. 3; for the XXZ model, we set $J=+1$, $B_{Z}=0.2$ and vary $J_{Z}\in[0,3]$ using truncated ASP with parameters $dt=0.1$ and $T_{\mathrm{max}}=1.2$. Note that for the XXZ model ASP is easier to implement due to the absence of protected crossings; the $B_{Z}$ field is introduced to break the (two-fold) degeneracy in the ground state.
@@
The low-energy subspace typically only contains a finite number of unique states. As we increase the number of training points, it is possible that we overspan the low-energy space, making the $\mathbb{S}$ matrix non-invertible. Thus, we only calculate errors for $k_{p}\leq(N_{XY}+2)$ for the XY model, and $k_{p}\leq N_{XXZ}$ for the XXZ model, which was empirically determined based on when the subspace basis vectors become degenerate.
@@
For both of these models, we observe that the number of basis vectors required for $(\Delta\varepsilon_{\mathrm{rel}})_{rms}<5\%$ scales linearly with the system size $N$, in stark contrast to the exponential scaling of the Hilbert space dimension. For the XY model, this is because the number of phase transitions (or level crossings) in the ground energy spectrum is proportional to $N$ (see Section D.1). For the XXZ model, since the ground state changes smoothly without any level crossings (Fig. A5), we required an even smaller number of training states, that again appear to scale linearly with system size ($N$).
@@ IV Avoiding protected level crossings and vanishing energy gaps
One of the benefits of using EC as a subspace method is that the subspace basis can be generated using Hamiltonians at points in parameter space that are easier to handle than the desired target parameters where state preparation methods perform poorly. For instance, ASP fails when the spectrum exhibits protected level crossings. In the previous section, these were broken by a small $B_{X}$ field; however, it may be that the $B_{X}=0$ problem is of interest. ITE can also have problems, and in particular, has issues with small energy gaps between the ground and excited states. In this section, we will illustrate how to ameliorate this issue using EC.
@@ IV.1 Getting around protected level crossings in ASP
Figure 5: Ground energy spectrum for XY model 5-site with parameters $\{J=1,B_{X}=0.0\}$ calculated using EC with subspace obtained from ASP in parameter space with $\{J=1,B_{X}=0.1$}
As discussed in Sec. III.0.2, using ASP to generate a ground state requires finding a gapped adiabatic path, that is without any protected crossings [24], which precludes its use for systems that exhibit level crossings. The 1D transverse field XY model is such a system: in the absence of an in-plane magnetic field — the spectrum shows several level crossings equal to the number of sites — and thus ASP is not an appropriate tool for finding the ground state. However, we can get around this issue by performing ASP for the XY model that does have a small in-plane field, and build a subspace using training states from this model.
In the Fig. 5, we perform ASP for a system with staggered $B_{X}=0.1$, $dt=0.05$ and $T_{\mathrm{max}}=3.75$, and choose four equally spaced states to form a basis. Next, we use this basis to find the eigenstates of the target system with $B_{X}=0.0$. As is clear from the figure, this procedure readily produces the correct ground state energies for the system of interest.
@@ V Noise resilience
In practice, since the currently available quantum hardware is noisy, we require algorithms to be noise resilient in order to yield useful results on a quantum device [11, 12, 13]. We already have some “noise” in the subspace in the form of excited states, arising from truncated state preparation methods; however, quantum devices introduce hardware noise as well, requiring us to analyze how additional noise will affect the algorithm. In this section, we will demonstrate that the improvements obtained using EC survive the addition of hardware noise. We implement a noise model and simulate the effect of introducing noise in the ASP for a 5-site XY model Hamiltonian discussed previously in Eq. (III.1).
@@
While increasing the number of training vectors (or subspace basis states) can help to reduce the error, this eventually reaches a barrier. Adding more basis vectors increases the chances of over-spanning the subspace, making the condition number $(C_{n})$ of the overlap matrix very large, and yielding it non-invertible (see Section II). This is one of the major problems with subspace methods ([33, 35, 69]). To address this issue, there are several classical post-processing techniques used to manipulate the overlap matrix, such as thresholding [43]. This procedure involves finding the singular value decomposition of the overlap matrix, followed by eliminating the singular values that are smaller than a chosen threshold ($\alpha$). The vectors corresponding to the singular values less than $\alpha$ are discarded from the overlap matrix and the projected Hamiltonian matrix. This technique can help in reducing the condition number but sometimes we can lose some relevant information in the process. Let us analyze the effect of noise on condition number and thresholding for our test case.
@@ VI Discussion
In this work, we have used previously proposed eigenvector continuation as a subspace method [48], where continuity of the ground state of the system Hamiltonians varied by some order parameter is explored. In contrast to previous work, where exact eigenstates were needed to generate the subspace, we apply it to the subspace generated by truncated time evolution techniques (ASP and ITE) and the variational quantum eigensolver (VQE) to generate a subspace. These truncated algorithms incorporate a number of low-energy states, which can be disentangled using EC, thus gaining better convergence while using fewer resources.
@@
We used various spin Hamiltonians with different ground state energy spectrum features to showcase the method, including the XY model with and without transverse field whose ground state manifold spans orthogonal symmetry sectors, the XXZ 1D model with a transverse field which has small energy gaps between ground state and the first excited state, and the XXZ Kagome 2D model which has phase transitions and ground state degeneracies. However, this method is not limited to spin systems; the eigenvector continuation formalism has been previously used for other systems in nuclear physics, and quantum chemistry among others [50, 52, 49], and the truncated state preparation method proposed in this work can also be used for these systems.
All the findings shown in this work were performed on a classical simulator. Even here, compared to the classical full-time-evolution methods to get the ground states, using EC with truncated time evolution can be much faster. For quantum simulations, however, it can be challenging to obtain the matrix elements of the projected Hamiltonian $\mathbb{H}$ and overlap matrix $\mathbb{S}$ [Eq. (II)]. One can evaluate these elements by performing multi-qubit controlled operations to an ancilla, e.g., the Hadamard test method [70]. These operations are known to be troublesome for current QPUs, as they introduce significant error. Other approaches for measuring overlap terms have been recently proposed to avoid the need to use the traditional Hadamard test, including a cheaper implementation when using ASP for producing the basis states [71], a SWAP test that could be cheaper than the Hadamard test [72], and an approach that relies on a reference state[33]. Moreover, algorithms for calculating the overlap of two quantum states efficiently on a quantum device remains an open research area.
While the calculation of the projected Hamiltonian and subspace matrices can involve otherwise expensive calculations, the number of these is limited. Because we use the same basis for all the target Hamiltonians, the overlap matrix elements need to be calculated once. Similarly, for the projected Hamiltonian matrix, usually, we do not need to repeat the measurements for each target Hamiltonian, since all the Pauli operators of the Hamiltonian are measured only once per basis set and only the coefficients are changed to calculate the target Hamiltonian [48]. Due to the structure of the matrices, the total number of circuits required scales quadratically in the number of basis states $k_{p}$: $k_{p}(k_{p}-1)$ for the overlap matrix and $k_{p}(k_{p}+1)$ for each of the Pauli operators in the projected Hamiltonian.
@@ Acknowledgements
This work was supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency. The variational quantum algorithm results were supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Award Number DE-SC0025623.
@@ D.1 XY model
Figure A3: XY 1D chain system energy spectrum for $J=1$, varying $B_{Z}$ and staggered $B_{X}\in\{0.0,0.2\}$. Level crossing split can be observed in the plot on the right.
Without a staggered X- field, the ground state is protected for a range of varying $B_{Z}$ due to the presence of magnetization symmetry sectors. For particular values of $B_{Z}$, the ground state changes to some excited state resulting in a point of degeneracy where both states cross called level crossing. The ground state before and after these points is completely orthogonal. For a $N$ site 1D chain system, for each spin flip, we will have a magnetization phase transition making the number of symmetry sectors in 1D transverse field XY model with varying transverse field from $B_{Z}\in(-\infty,+\infty)$ as $(N+1)$. For example we plot $N=8$ energy spectrum with $J=1$, staggered $B_{X}=0.0$ and $B_{Z}\in(-3,3)$ in Fig. A4. For the plots in this paper, we consider $B_{Z}\in[0,B_{Z(max)})$. Still, the scaling of the number of symmetry sectors would be linear in $N$.
@@
Adding a uniform field in the X-direction of strength $B_{X}$ has two significant consequences. One is the breaking of level crossings, and the other is the mixing of the lowest eigenstates. Breaking of level crossing introduces an energy gap between the lowest energy states, which makes the application of ASP possible. After adding a staggered $B_{X}$ field, the ground states are not completely orthogonal unlike one without the field. This makes the training states mixed as well.
@@ D.2 XXZ model 1D chain
Figure A6: Implementing EC with truncated ASP and ITE for a 1D chain XXZ model 5 site. (a) and (c) plots the ground energy spectrum, (b) and (d) plots corresponding ground state fidelity change.
In the Fig. A6, truncated ITE and ASP are used to get a basis using and for XXZ 1D chain Hamiltonians D.2, with $J=1$, $B_{Z}=0.2$, and $J_{Z}\in(0,3)$ is the varying parameter. For truncated ITE in (a) and (b), $d\tau=0.2$ and $\tau_{\mathrm{max}}=1.2$ starting from a Haar random state. And for ASP in (c) and (d), starting from a ground state at $J_{Z}=0$, $4$ time steps with $dt=0.1$ and $T_{\mathrm{max}}=0.4$ are used to obtain the basis.
Figure A7: Energy spectrum of a 12-site Kagome Lattice system with $N_{X}=N_{Y}=2$, $J=1.0$, $B_{Z}=0.2$ and varying $J_{Z}$ showing lowest $100$ energy levels
@@ Appendix E Noise analysis
E.1 Measurement error
To account for the measurement errors in the noise analysis, we introduce an error to all the matrix element values. Measurement errors could include two main types of errors including a readout error caused by an error in reading the output and another due to $T1$ decoherence error causing bit-flip from $0\rightarrow 1$. As mentioned later in the discussion, there could be multiple ways of calculating the matrix elements on a quantum device. For our measurement error analysis, we considered the Hadamard test protocol. To do so, we consider all the samples generated by implementing the noise model and change the expected matrix element values according to the following equation. We consider the values of readout errors from a ref [73] as $p_{0\rightarrow 1}=0.01$ and $p_{1\rightarrow 0}=0.03$. To account for $T_{1}$ error decoherence, we add a max $p_{T1}=0.03$ that cause a biased bit-flip from |$\ket{1}\rightarrow\ket{0}$.
For comparison with ASP, we included measurement errors in measuring Hamiltonian expectation values for a state produced by an ASP circuit by introducing a confusion matrix as explored by Rigetti in [74]. In the A10, the shaded region shows values with measurement error for ASP and with EC.
Figure A10: Noise analysis with measurement error. The shaded region shows the change in error including the measurement error.